# BOOST: Bayesian Optimization with Optimal Kernel and Acquisition Function Selection Technique

### v1

@@ Methods
  • Acquisition functions: {EI, PI, UCB ($\beta=0.1$), PM}
@@
The internal BO process is limited to 20 iterations for computational efficiency. This efficiency is further enhanced by the fact that the data-based evaluation in BOOST is fully parallelizable, as each kernel–acquisition pair is assessed independently.
Recommendation
@@
  • Acquisition functions: EI $>$ PI $>$ UCB $>$ PM
  • Kernel functions: Matérn 3/2 $>$ Matérn 5/2 $>$ RBF $>$ RQ
This priority is based on the degree of exploration each acquisition function promotes. When data is scarce, exploration is essential to acquire informative samples that help reveal the structural characteristics of the objective function. This, in turn, enables the internal BO process to more effectively evaluate and distinguish between different kernel–acquisition combinations. Accordingly, we prioritize acquisition functions with stronger exploratory tendencies, ranking them as EI $>$ PI $>$ UCB $>$ PM. In particular, UCB is placed below PI because we adopt a low exploration coefficient $(\beta=0.1)$, which makes it behave almost greedily (Papenmeier et al. 2025; Dogan and Prestwich 2022; Tian et al. 2024).
Similarly, we prioritize Matérn kernels over RBF due to their greater flexibility in modeling non-smooth or complex objective functions, as they are recommended for many realistic problems where strong smoothness assumptions are unrealistic (Snoek et al. 2012; Williams and Rasmussen 2006; Stein 1999). We place RBF over RQ since RBF is the most widely used kernel, while RQ is a special case of kernels which can be written as a mixture of RBF kernels with various length scales (Williams and Rasmussen 2006).
@@
BOOST 1 2 1 2 2 1 2 3 1 1.67
Matérn32_EI 10 1 3 4 4 2 4 6 9 4.78
Matérn52_EI 5 4 2 3 11 12 7 5 4 5.89
Matérn52_PI 2 10 8 9 5 5 3 2 16 6.67
RQ_EI 12 3 7 1 3 15 6 8 8 7.00
Matérn32_PI 3 13 6 16 10 9 9 4 10 8.89
RBF_PI 17 7 17 5 15 17 1 1 2 9.11
RQ_PI 9 6 14 10 14 3 8 14 5 9.22
RBF_EI 16 5 12 7 1 16 5 7 17 9.56
RBF_UCB 13 8 11 6 13 11 12 15 6 10.56
RQ_PM 8 11 15 13 8 6 10 9 15 10.56
Matérn32_PM 14 16 10 17 7 8 13 10 3 10.89
Matérn32_UCB 15 14 9 15 6 4 15 16 7 11.22
RQ_UCB 7 12 13 12 9 14 11 13 11 11.33
Matérn52_UCB 4 15 4 11 16 13 16 12 12 11.44
Matérn52_PM 6 17 5 14 12 10 17 11 13 11.67
RBF_PM 11 9 16 8 17 7 14 17 14 12.56
Table 1: Final performance rankings (lower is better) based on regret after 100 samples across four synthetic functions and five HPO-B benchmarks. Numbers 5636–5906 indicate the search space IDs used in the HPO-B dataset. To ensure fair ranking, if the regrets after $n$ evaluations were tied, we compared regrets at $n-1$ to resolve ties.
Experiments
We conduct comprehensive experiments to evaluate BOOST’s performance across various discrete optimization scenarios. All benchmark tasks are designed in fully discretized spaces to align with practical problem settings and to demonstrate BOOST’s real-world applicability (González-Duque et al. 2024). We compare BOOST against fixed kernel–acquisition function combinations on both synthetic benchmark functions and real-world hyperparameter optimization tasks.
@@
Synthetic Benchmark Functions
We evaluate BOOST on four widely used 4-dimensional discrete synthetic benchmark functions: Ackley, Levy, Rosenbrock, and SumSquares (Surjanovic and Bingham 2025). Detailed discretization parameters are provided in the Appendix. We use Latin Hypercube Sampling (LHS) to select 10 initial points, ensuring well-distributed coverage of the search space.
Real-world machine learning hyperparameter optimization tasks
To assess the effectiveness of BOOST in real-world problems, we conduct experiments on five hyperparameter optimization tasks selected from the HPO-B dataset (Arango et al. 2021). These tasks involve Decision Trees (6D), Random Forests (9D), and XGBoost (16D) with varying search space dimensionalities. For each task, we randomly select 10 initial points from the available dataset, as the search space is predefined and sampling methods such as LHS are not applicable. Detailed experimental configurations and data preprocessing steps are provided in the Appendix.
Experimental Setup
We compare BOOST against 16 deterministic baseline methods, representing all possible combinations of the four kernel functions (Matérn 3/2, Matérn 5/2, RBF, RQ) and four acquisition functions (EI, PI, UCB, PM) used in our candidate set.
Results
As shown in Figure 2, BOOST consistently demonstrates competitive performance across both synthetic benchmark functions and real-world machine learning hyperparameter optimization tasks. Compared to other fixed hyperparameter methods, BOOST ranked within the top 3 across all test cases when comparing regret after 100 evaluations (10 initial samples followed by 90 BO iterations), highlighting its adaptability and effectiveness in diverse optimization scenarios. To ensure fair ranking, if the regrets after $n$ evaluations were tied, we compared regrets at $n-1$ to resolve ties.
BOOST achieved the best overall performance with an average rank of 1.67 out of 18 candidates (Table 1). The second-best method, Matérn 3/2 with EI, had an average rank of 4.78. This suggests that BOOST is not only competitive on simple problems but also robust across both simple and complex functions, consistently identifying suitable kernel and acquisition combinations for any possible search space.
Conclusion
@@
Unlike fixed hyperparameter methods that may excel in specific settings but fail in others, BOOST dynamically adapts its strategy based on observed data characteristics, eliminating the need for manual tuning while ensuring consistent performance across diverse optimization scenarios. Our results demonstrate that adaptive hyperparameter selection can achieve notable performance gains compared to fixed hyperparameter methods.
@@ Appendix A Search Space Description
Discretized Synthetic Benchmark Functions
@@
Ackley $f(x)=-a\exp\left(-b\sqrt{\frac{1}{d}\sum x_{i}^{2}}\right)-\exp\left(\frac{1}{d}\sum\cos(cx_{i})\right)+a+e$, where $a=20,\ b=0.2,\ c=2\pi$.
@@
where $w_{i}=1+\frac{x_{i}-1}{4}$.
Rosenbrock $f(x)=\sum_{i=1}^{d-1}[100(x_{i+1}-x_{i}^{2})^{2}+(1-x_{i})^{2}]$.
@@
Machine Learning Hyperparameter Optimization Tasks
@@ Appendix B Information for Reproducibility
All experiments were repeated 10 times to ensure statistical reliability. To ensure reproducibility, we controlled randomness by adopting a consistent seeding strategy: for the $n$th trial, we used seed $n-1$ (i.e., seeds 0–9). For K-means clustering, we fixed the random seed to 42 for consistency with common practice
Hyperparameter Settings
All key hyperparameters used in our experiments are listed below. Additional configuration details, such as the number of initial samples and optimization iterations, are provided in the main text.
@@
    • Noise constraint: $[5\times 10^{-4},0.2]$
    • Length scale constraint: $[5\times 10^{-6},\sqrt{d}]$, where $d$ is the input dimension
    • Output scale constraint: [0.05,20.0][0.05,20.0][ 0.05 , 20.0 ]
@@
  • pandas: 2.2.3
Appendix C Additional results
@@
Effect of Data Partitioning Method
@@
Effect of Data Partitioning Size
@@
Effect of Stopping Criteria
@@
Effect of Tie-Breaking Rule

### v2

@@ 4.3 Performance Evaluation
  • Acquisition functions: {EI, PI, LCB ($\beta=0.1$), PM}
@@
The data-based evaluation in BOOST is fully parallelizable, as each kernel–acquisition pair is assessed independently. In expensive black-box optimization where function evaluations can take hours to days, what truly matters is the efficiency and accuracy of the proposed evaluation points, rather than marginal computational overhead on the order of seconds (Snoek et al.,, 2012; Cosenza et al.,, 2022). Detailed runtime analysis is provided in  D.
4.4 Recommendation
@@
  • Acquisition functions: EI >> PI >> LCB >> PM
  • Kernel functions: Matérn 3/2 >> Matérn 5/2 >> RBF >> RQ
This priority is based on the degree of exploration each acquisition function promotes. When data is scarce, exploration is essential to acquire informative samples that help reveal the structural characteristics of the objective function. This, in turn, enables the internal BO process to more effectively evaluate and distinguish between different kernel–acquisition combinations. Accordingly, we prioritize acquisition functions with stronger exploratory tendencies, ranking them as EI >> PI >> LCB >> PM. In particular, LCB is placed below PI because we adopt a low exploration coefficient $(\beta=0.1)$, which makes it behave almost greedily (Dogan and Prestwich,, 2022; Tian et al.,, 2024; Papenmeier et al.,, 2025).
Similarly, we prioritize Matérn kernels over RBF due to their greater flexibility in modeling non-smooth or complex objective functions, as they are recommended for many realistic problems where strong smoothness assumptions are unrealistic (Stein,, 1999; Williams and Rasmussen,, 2006; Snoek et al.,, 2012). We place RBF over RQ since RBF is the most widely used kernel, while RQ is a special case of kernels which can be written as a mixture of RBF kernels with various length scales (Williams and Rasmussen,, 2006).
@@
Figure 3: Optimization performance comparison between BOOST and state-of-the-art/adaptive methods on the same benchmark tasks as Figure 2.
5 Experiments
We conduct comprehensive experiments to evaluate BOOST’s performance across various discrete optimization scenarios. All benchmark tasks are designed in fully discretized spaces to align with practical problem settings and to demonstrate BOOST’s real-world applicability (González-Duque et al.,, 2024). We compare BOOST against fixed kernel–acquisition function combinations, a state-of-the-art method, and adaptive approaches on both synthetic benchmark functions and real-world hyperparameter optimization tasks.
@@ 5 Experiments
5.1 Synthetic Benchmark Functions
We evaluate BOOST on four widely used 4-dimensional discrete synthetic benchmark functions: Ackley, Levy, Rosenbrock, and SumSquares (Surjanovic and Bingham,, 2025). Detailed discretization parameters are provided in Appendix A.1. We use Latin Hypercube Sampling (LHS) to select 10 initial points, ensuring well-distributed coverage of the search space.
5.2 Real-world machine learning hyperparameter optimization tasks
To assess the effectiveness of BOOST in real-world problems, we conduct experiments on five hyperparameter optimization tasks selected from the HPO-B dataset (Arango et al.,, 2021). These tasks involve Decision Trees (6D), Random Forests (9D), and XGBoost (16D) with varying search space dimensionalities. For each task, we randomly select 10 initial points from the available dataset, as the search space is predefined and sampling methods such as LHS are not applicable. Detailed experimental configurations and data preprocessing steps are provided in Appendix A.2.
5.3 Experimental Setup
We first compare BOOST against 16 deterministic baseline methods, representing all possible combinations of the four kernel functions (Matérn 3/2, Matérn 5/2, RBF, RQ) and four acquisition functions (EI, PI, LCB, PM) used in our candidate set.
We then compare BOOST with a state-of-the-art algorithm and adaptive approaches: SAASBO (Eriksson and Jankowiak,, 2021), a widely recognized state-of-the-art algorithm; Best Utility (Roman et al.,, 2019), an adaptive kernel method; and No-PASt-BO (Vasconcelos et al.,, 2019), an adaptive acquisition method.
5.4 Results
As shown in Figure 2 and Figure 3, BOOST consistently demonstrates competitive performance across both synthetic benchmark functions and real-world machine learning hyperparameter optimization tasks. Compared to other fixed hyperparameter methods (Figure 2), BOOST ranked within the top 3 across all test cases when comparing regret after 100 evaluations (10 initial samples followed by 90 BO iterations), achieving the best overall performance with an average rank of 1.67 out of 17 candidates (Table 3 in Appendix E). The second-best method, Matérn 3/2 with EI, had an average rank of 4.78, highlighting BOOST’s significant advantage. To ensure fair ranking, if the regrets after $n$ evaluations were tied, we compared regrets at n1n-1 to resolve ties.
Even compared with state-of-the-art and adaptive approaches (Figure 3), BOOST maintained its superiority. Except for two simple problems—Rosenbrock (valley-shaped) and SumSquares (bowl-shaped)—BOOST either dominated or remained highly competitive across all test cases. This demonstrates BOOST’s particular strength in complex real-world optimization scenarios.
6 Conclusion
@@
Unlike fixed hyperparameter methods that may excel in specific settings but fail in others, BOOST dynamically adapts its strategy based on observed data characteristics, eliminating the need for manual tuning while ensuring consistent performance across diverse optimization scenarios. Our results demonstrate that adaptive hyperparameter selection can achieve notable performance gains even compared to state-of-the-art algorithms.
Appendices
@@ Appendix A Search Space Description
A.1 Discretized Synthetic Benchmark Functions
@@
Ackley $f(x)=-a\exp\!\left(-b\sqrt{\tfrac{1}{d}\sum x_{i}^{2}}\right)-\exp\!\left(\tfrac{1}{d}\sum\cos(cx_{i})\right)+a+e$,
where $a=20,\ b=0.2,\ c=2\pi$.
@@
where $w_{i}=1+\tfrac{x_{i}-1}{4}$.
Rosenbrock $f(x)=\sum_{i=1}^{d-1}\left[100(x_{i+1}-x_{i}^{2})^{2}+(1-x_{i})^{2}\right]$.
@@
A.2 Machine Learning Hyperparameter Optimization Tasks
@@ Appendix B Information for Reproducibility
All experiments were repeated 10 times to ensure statistical reliability. To ensure reproducibility, we controlled randomness by adopting a consistent seeding strategy: for the $n$th trial, we used seed n1n-1 (i.e., seeds 0–9). For K-means clustering, we fixed the random seed to 42 for consistency with common practice
B.1 Hyperparameter Settings
All key hyperparameters used in our experiments are listed below. Additional configuration details, such as the number of initial samples and optimization iterations, are provided in the main text. For the baselines SAASBO and No-PASt-BO, we used the default settings provided in their official code without modification. For Best Utility, since the code was not publicly available, we implemented it ourselves based on the description in the original paper, using the same settings as BOOST to ensure a fair comparison.
@@ B.1 Hyperparameter Settings
    • Noise constraint: $5\times 10^{-4},0.2]$
    • Length scale constraint: $5\times 10^{-6},\sqrt{d}]$, where $d$ is the input dimension
    • Output scale constraint: $[0.05,20.0]$
@@ B.3 Software
For the No-PASt-BO baseline, the official code relies on the GPyOpt package, which is not compatible with recent versions. Therefore, we ran the baseline in a separate environment with older versions of NumPy/SciPy in a separate environment:
  • Python: 3.11.11
  • PyTorch: 2.7.1
  • torchvision: 0.22.1
  • torchaudio: 2.7.1
  • GPy: 1.13.2
  • GPyOpt: 1.2.6
  • NumPy: 1.26.4
  • SciPy: 1.12.0
  • scikit-learn: 1.6.1
Appendix C Pseudocode for BOOST
Algorithm 1 Overall algorithm for BOOST
Output: Best Kernel-Acquitision function pair ($k^{*}$, $\alpha^{*}$)
[…]
5:for all $(k_{i},\alpha_{j})$ in $\mathcal{K}\times\mathcal{A}$ do
6:  $R_{0}\leftarrow r_{n}=\{(x_{i},f(x_{i}))\}_{i=1}^{n_{init}}$
7:  $Q_{0}\leftarrow\{x\mid(x,f(x))\in q_{n}\}$
8:  $t\leftarrow 0$
9:  while $t<T_{\max}$ do
10:   Train a Gaussian Process using $k_{i}$ over $R_{t}$
11:   $x_{t+1}\leftarrow\arg\max_{x\in Q_{t}}\alpha_{j}(x)$
12:   $R_{t+1}\leftarrow R_{t}\cup\{(x_{t+1},f(x_{t+1}))\}$
13:   $Q_{t+1}\leftarrow Q_{t}\setminus\{x_{t+1}\}$
14:   $t\leftarrow t+1$
15:   if $f(x_{t+1})\leq y_{target}$ then
16:    break
17:   end if
18:  end while
19:  $t_{k_{i},\alpha_{j}}\leftarrow t$
20:end for
Appendix D Time Analysis
We compared the total runtime for 90 BO iterations on the Levy function. As described in the main text, all experiments used 10 initial samples followed by 90 optimization iterations (100 evaluations in total). The experiments were conducted on an AMD Ryzen 9 7900X CPU with 12 cores, allowing extensive parallelization across algorithms. For BOOST, although 16 candidate kernel–acquisition pairs were evaluated at each iteration, we limited the number of parallel workers to 8 (out of 12 available cores) to balance runtime efficiency and resource usage. For the other algorithms, parallel computation was applied whenever the method naturally allowed simultaneous evaluations. To preserve fidelity to the original algorithms, the code and data appendices contain results from the sequential implementations. However, in this section, we report results obtained by modifying the code to exploit parallelization for a fair comparison of runtime.
Table 2: Total runtime (90 iterations) and average runtime per iteration on the Levy function.
Algorithm Total Runtime (s) Avg. per Iter. (s)
BOOST (Ours) 608.27 6.76
Matérn32 EI 163.05 1.81
SAASBO 1977.08 21.97
Best Utility 569.34 6.32
No-PASt-BO 536.60 5.96
Appendix E Additional Analysis of Results
Table 3: Final performance rankings (lower is better) based on regret after 100 samples across four synthetic functions and five HPO-B benchmarks. Numbers 5636–5906 indicate the search space IDs used in the HPO-B dataset. To ensure fair ranking, if the regrets after $n$ evaluations were tied, we compared regrets at n1n-1 to resolve ties.
BOOST (Ours) 1 2 1 2 2 1 2 3 1 1.67
Matérn32_EI 10 1 3 4 4 2 4 6 9 4.78
Matérn52_EI 5 4 2 3 11 12 7 5 4 5.89
Matérn52_PI 2 10 8 9 5 5 3 2 16 6.67
RQ_EI 12 3 7 1 3 15 6 8 8 7.00
Matérn32_PI 3 13 6 16 10 9 9 4 10 8.89
RBF_PI 17 7 17 5 15 17 1 1 2 9.11
RQ_PI 9 6 14 10 14 3 8 14 5 9.22
RBF_EI 16 5 12 7 1 16 5 7 17 9.56
RBF_LCB 13 8 11 6 13 11 12 15 6 10.56
RQ_PM 8 11 15 13 8 6 10 9 15 10.56
Matérn32_PM 14 16 10 17 7 8 13 10 3 10.89
Matérn32_LCB 15 14 9 15 6 4 15 16 7 11.22
RQ_LCB 7 12 13 12 9 14 11 13 11 11.33
Matérn52_LCB 4 15 4 11 16 13 16 12 12 11.44
Matérn52_PM 6 17 5 14 12 10 17 11 13 11.67
RBF_PM 11 9 16 8 17 7 14 17 14 12.56
Appendix F Additional Experiments and Results
@@
F.1 Effect of Data Partitioning Method
@@
F.2 Effect of Data Partitioning Size
@@
Importantly, this robustness across different partitioning ratios demonstrates the effectiveness of BOOST’s core principle: evaluating configurations based on retrospective performance within the data-in-hand provides a reliable and stable selection mechanism. The consistent performance across various ratios suggests that our approach successfully captures the inherent characteristics of different kernel-acquisition pairs, making it less sensitive to specific partitioning sizes.
@@
F.3 Effect of Stopping Criteria
@@
F.4 Effect of Tie-Breaking Rule